OpenStudy (aripotta):

If m and k are positive and 10m^2k^-1 = 100m, what is m^-1 in terms of k? a. k/10 b. k/90 c. sqrt(k)/10 d. 1/10k e. 1/90k

OpenStudy (aripotta):

\[10m ^{2}k ^{-1} = 100m\]

ganeshie8 (ganeshie8):

dividing both sides by \(m^2\) gives you \(m^{-1}\) on right hand side

ganeshie8 (ganeshie8):

if that doesn't look obvious, how about solving the equation for \(m\) ?

OpenStudy (aripotta):

divide by 100?

ganeshie8 (ganeshie8):

yeah but there is m on both sides right ?

OpenStudy (aripotta):

yea

ganeshie8 (ganeshie8):

maybe divide \(m^2\) both sides and see what u get

ganeshie8 (ganeshie8):

\[\large \rm 10m ^{2}k ^{-1} = 100m\] dividing \(\large \rm m^2\) both sides gives \[\large \rm \dfrac{10m ^{2}k ^{-1}}{m^2} = \dfrac{100m}{m^2}\]

OpenStudy (aripotta):

ok yea i was trying to type that ^

ganeshie8 (ganeshie8):

next divide 100 both sides so that m will be isolated all by itself

OpenStudy (aripotta):

why divide?

ganeshie8 (ganeshie8):

becaule you want to isolate m, and since 100 is attached to m, you want to get rid of it

ganeshie8 (ganeshie8):

dividing 100 both sides kills the 100 on right hand side

ganeshie8 (ganeshie8):

\[\large \rm 10k ^{-1} = \dfrac{100}{m}\] dividing 100 both sides you get \[\large \rm \dfrac{10k ^{-1}}{100} = \dfrac{1}{m}\]

ganeshie8 (ganeshie8):

\[\large \rm \dfrac{k ^{-1}}{10} = \dfrac{1}{m}\]

OpenStudy (aripotta):

which is equal to 1/10k right

ganeshie8 (ganeshie8):

Yep ! notice the right hand side is same as \(\large \rm m^{-1}\)

OpenStudy (aripotta):

alrighty, thank you a lot. it started making sense once you started working it out. i felt stupid for not realizing it before i posted it